# references.bib

### v1

@@ 3.2 A non-linear propagator model
The above results can be summarized within the framework of a non-linear propagator model [bouchaud2003fluctuations, bouchaud2018trades], where the impact of child order $j$ measured at time $t_{i}>t_{j}$ is proportional to $\sqrt{q_{j}/(t_{i}-t_{j}+s_{0})}$, as predicted both by the LLOB model [donier2015fullyconsistentminimalmodel] and by the Bayesian theory of Ref. [saddier2024bayesian]. Indeed, from such an expression one gets:
@@
Such an interpretation however appears to violate the “diffusivity” condition derived in [bouchaud2003fluctuations, jusselin2020no], which relates the decay of the propagator with the decay of the autocorrelation of the sign of the trades. Superficially, a propagator decay in $(t_{i}-t_{j})^{-1/2}$ should lead to strongly mean reverting prices, at odds with the diffusive nature of prices. A way out of this conundrum will be presented in a forthcoming paper [ustocome]. Note that the above non-linear propagator model precisely saturates the no-arbitrage bound derived by Gatheral in [gatheral2010no].
The most striking result of the previous Section is that the square-root law appears to hold already at the level of child orders. This is not in line with the standard “mesoscopic” interpretation of the Latent Liquidity model [donier2015fullyconsistentminimalmodel], which, as we alluded to before, would require the fraction of slow volume to be much smaller than what we reported in Section 2.3.
We are thus led to the conclusion that the latent liquidity idea must in fact operate already at the micro level. Whereas the revealed order book contains primarily limit orders posted by market makers, the final sellers’ or buyers’ price would be distributed according to a locally linear profile, as predicted by the LLOB theory [donier2015fullyconsistentminimalmodel] – which, as a reminder, only relies on minimal assumptions, in particular on the diffusive nature of prices.
So the scenario would be as follows: once the incoming buy (sell) has been executed against a market-maker, a “hot potato” game starts between market-makers until the order is finally digested by a final seller (buyer). In order to find such a final seller (buyer) the price must on average move by an amount $\delta p$ such that $\Gamma(\delta p)^{2}/2=q$, where $\Gamma$ is the slope of the latent liquidity. Once this is achieved, the price tends to revert back as $1/\sqrt{t-t_{i}}$, as predicted by the LLOB model, see [donier2015fullyconsistentminimalmodel, bouchaud2018trades], but in disagreement with the predictions of the propagator model [bouchaud2003fluctuations]. This discrepancy will be discussed further in [ustocome].
If the above interpretation is correct, however, it should hold for arbitrary market orders. Since all orders are equivalent, they should lead to the same average impact on prices (as was indeed found in [toth2017short]). In other words, one should see a $\sqrt{q}$ impact for single market orders provided one waits long enough for the “hot potato” game to be completed. This is what we test in the next Section, which then opens up the question of reconstructing synthetic metaorders from a list of consecutive market orders that do not necessarily belong to the same ID. The anonymity of market orders suggests that, in certain conditions, the very same square-root impact law $\mathcal{I}$ given by Eq. (1) should also hold for synthetic metaorders. This indeed turns out to be the case, as we discuss now.
@@ 4.1 The impact of single public market orders
We want to understand the behavior of the price after a buy (sell) market order. Clearly, if the volume of the market order is less than the prevailing volume at the opposite best, the immediate impact is zero. However, as time goes by, one very quickly sees that impact grows and becomes approximately given by $\sqrt{q}$, whether or not immediate impact is zero (the full temporal aspects will be explored in more details in [ustocome]).
@@ 4.2 Synthetic metaorders
Since arbitrary market orders seem to all behave similarly, independently of the metaorder they belong to, we made the following numerical experiment that confirms the non-linear propagator interpretation of Eq. (2) for metaorders. We construct a new dataset of synthetic metaorders, by randomly shuffling traders ID and distributing them to market orders while we keep the historical market order flow, i.e. by randomly reordering real traders’ IDs. This preserves the initial frequency distribution of traders: i.e. some of them appear many times whereas others are trading less frequently.
We then use the same method as in 2.1 to define metaorders as a consecutive sequence of trades of the same sign associated to the same new trader ID (that has been reshuffled). We obtain synthetic metaorders, that start and end at different times as the original (true) metaorders. Hence, any information associated with these metaorders is at least partially lost. Still, as shown in Fig. 6, we recover exactly the same square-root impact function as for the original metaorders! Note in passing that these graphs show once again that the square-root impact law only depends on the volume $Q$ of the metaorder (either real or synthetic) but not of the execution time $T$, see [bucci_impact_vol2019, bouchaud2018trades, kyle2018market]. We have tested different constructions of synthetic metaorders on different stocks, from the Paris and London Stock Exchange, with similar results – a more detailed discussion will be presented in [ustocome].
Figure 6: Left: Impact $\mathcal{I}(Q)$ of real metaorders as a function of their rescaled size $Q/V_{D}$ for a typical Japanese Stock, using data from 2012 to 2018. The color of the dots correspond to different total execution time $T$, expressed in seconds. Right: Impact $\mathcal{I}(Q)$ of synthetic metaorders for the same stock, obtained via ID reshuffling. ID shuffling consists in a random permutation of historical trader IDs, preserving the frequency of apparition. Note that the vertical and horizontal scales are the same in the two plots: the square-root fit is exactly the same as for real metaorders. From the legend, one can also clearly see that $\mathcal{I}(Q)$ is independent of $T$ [bucci_impact_vol2019].
@@ 4.3 Discussion
The results of the previous two sub-sections strongly suggest a purely “mechanical” interpretation of the square-root impact law, based on a time decaying $\sqrt{q/t}$ impact of single market orders, independently of their association with a specific metaorder (since trader ID’s can be scrambled without affecting the results). These findings are difficult to reconcile with theories explaining the square-root law based on information, such as in [gabaix2006], or on the detection by the market of the beginning of new metaorders, such as in [farmer2013efficiency, saddier2024bayesian].
Those results are, on the other hand, perfectly in line with the fact that, due to anonymity, all market orders – even uninformed ones – should play an equivalent role and should on average impact prices similarly [bouchaud2010impact]. This was already noted in [toth2017short] by comparing the impact of CFM market orders with non-CFM market orders, and even more convincingly in an unpublished specifically designed 2010 experimental campaign with totally random market orders.
@@ 5.1 Refill sequences
Whereas the long-term correlation of market orders is well documented [bouchaud2003fluctuations, lillo2005theory, bouchaud2009markets] and mainly attributed to the order splitting of large metaorders [lillo2005theory, sato2023inferring], our dataset also allows us to study the splitting strategy of liquidity providers.
To do so, we simplify the problem by restricting to the set of filled limit orders, i.e. limit orders that have been placed in the order book and subsequently executed by another participant. Thus, as for liquidity takers, one can aggregate those filled limit orders into “refill sequences”, i.e. sequences of consecutive filled limit orders of same sign submitted by the same trader during a trading session. Given the splitting behavior of liquidity takers, market makers/liquidity providers face a sign-correlated succession of market orders. It is thus likely that the executed limit orders flow will be also be persistent, and lead to a power-law tail in the size distribution of the refill sequences, as we indeed confirm empirically, see Fig. 7. We show there a power-law fit of the distribution of the number $n$ of child orders associated with refill sequences, as $\psi(n)\propto n^{-\mu_{p}}$ with $\mu_{p}$ in the range $[1.4,2.4]$ depending on the considered stock. This power-law decay echoes the Lillo-Mike-Farmer distribution of child market orders, although it is not expected to mirror it exactly since different traders provide liquidity to the same incoming metaorder – see [eisler2012price] for a related discussion.
It should however be emphasized at this point that the boundary between liquidity provider and liquidity taker is somewhat blurred. Except for specific cases, the large majority of market participants use a mix of limit and market orders to acquire or sell shares. For example, it is quite common to see participants adding liquidity at the bid (ask) while sending market orders at the ask (bid). Large funds like AQR declare that most of their executed volume is though limit orders [AQR].
@@ 5.2 Strategic behaviour of liquidity providers
  • High $C_{\ell}\gtrsim 0.02$ traders are “wary” and place their next limit order quite a bit deeper in the book. The corresponding values of $\kappa_{\ell}$ cluster around $1/2$.
@@
Low $C_{\ell}$ market makers are thus responsible for ensuring stable liquidity. Figure 9 confirms that low $C_{\ell}$ traders account for the largest fraction of consumed liquidity. Note that $C_{\ell}=0.02$ corresponds to a price degradation of $2\%$ of the daily volatility $\sigma_{D}$ after the first executed limit order, and of $0.2\%$ after the 10th execution when $\kappa_{\ell}=1$.
@@ 6 Conclusion
The JPX database provides a trove of interesting features, which have only started to be exploited by Sato and Kanazawa to understand the origin of the long memory of market order signs [sato2023inferring] and, more recently, to firmly establish the square-root law of market impact, Eq. (1) and rule out some of the proposed theories that predict a non-universal value for the concavity exponent $\delta$ [kanazawasato2024]. Our aim in this study was to leverage the fact that all metaorders can be identified to shed light on the microscopic origin of the square-root law. Our central result, which we did not expect when starting this project, is that such a law has in fact microscopic roots, and applies already at the level of single child orders, provided one waits long enough for the market to “digest” these orders. This is not consistent with the standard interpretation of the LLOB model [donier2015fullyconsistentminimalmodel], which assumed that the theory described the liquidity dynamics at a “mesoscopic” scale. The mesoscopic impact of metaorders rather arises from a “double” square-root effect at the level of child orders, see Eq. (2): square-root in volume of individual impact, followed by an inverse square root decay as a function of child order time, such that the cumulative impact of a metaorder yields back the partial (Eq. (4)) and total (Eq. (3)) square-root laws.
@@
This finding however immediately suggests that since market orders are anonymous, the double square-root law Eq. (2) should apply to any market orders and the impact of synthetic metaorders, reconstructed by scrambling the identity of the issuers, should also be described by the square-root impact law, Eq. (1). We have provided empirical evidence that this both statements are indeed valid. In particular, synthetic metaorders behave exactly as real metaorders, see Fig. 6. We conclude that there is nothing special about child orders belonging to a given metaorder, at odds with theories that emphasize the information content of such trades to explain the square-root impact law, but in agreement with previous conjectures about the purely mechanical aspect of price impact [bouchaud2010impact, toth2017short, donier2016walras].
@@
Interestingly, our synthetic metaorder experiment suggests that it may be possible to reconstruct the impact of metaorders from the public tape only, without trader IDs. In a forthcoming publication [ustocome], we show that this is indeed the case, provided market orders are properly aggregated into synthetic metaorders, opening the path to a new wave of empirical studies, in particular concerning cross-impact [benzaquen2017dissecting, hey2023trading].
@@
While our results show that the square-root impact law does not emerge at the meso-scale but is already present at the micro-scale, they trigger new unanswered questions. In particular, why is the impact of single market orders of volume $q$ also a square-root? We have argued that this is because the latent order book is locally linear, such that after a “hot potato” game between liquidity providers, the final counterparty of the initial market order is on average at a distance $\sqrt{q}$ from the initial mid-point. Although this scenario is intuitively plausible, we believe that a deeper dive into the JPX database (or a similar one) would allow one to (in-)validate such a picture. Another conundrum is the square-root time decay of individual market orders, which seems to violate the martingale constraint that relates the decay of the autocorrelation of the sign of market orders to impact decay [bouchaud2003fluctuations, jusselin2020no]. We hope to come back to these pressing issues in the near future.

### v2

@@ 3.2 A non-linear propagator model
The above results can be summarized within the framework of a non-linear propagator model [16, 7], where the impact of child order $j$ measured at time $t_{i}>t_{j}$ is proportional to $\sqrt{q_{j}/(t_{i}-t_{j}+s_{0})}$, as predicted both by the LLOB model [22] and by the Bayesian theory of Ref. [21]. Indeed, from such an expression one gets:
@@
Such an interpretation however appears to violate the “diffusivity” condition derived in [16, 40], which relates the decay of the propagator with the decay of the autocorrelation of the sign of the trades. Superficially, a propagator decay in $(t_{i}-t_{j})^{-1/2}$ should lead to strongly mean reverting prices, at odds with the diffusive nature of prices. A way out of this conundrum will be presented in a forthcoming paper [41]. Note that the above non-linear propagator model precisely saturates the no-arbitrage bound derived by Gatheral in [42].
The most striking result of the previous Section is that the square-root law appears to hold already at the level of child orders. This is not in line with the standard “mesoscopic” interpretation of the Latent Liquidity model [22], which, as we alluded to before, would require the fraction of slow volume to be much smaller than what we reported in Section 2.3.
We are thus led to the conclusion that the latent liquidity idea must in fact operate already at the micro level. Whereas the revealed order book contains primarily limit orders posted by market makers, the final sellers’ or buyers’ price would be distributed according to a locally linear profile, as predicted by the LLOB theory [22] – which, as a reminder, only relies on minimal assumptions, in particular on the diffusive nature of prices.
So the scenario would be as follows: once the incoming buy (sell) has been executed against a market-maker, a “hot potato” game starts between market-makers until the order is finally digested by a final seller (buyer). In order to find such a final seller (buyer) the price must on average move by an amount $\delta p$ such that $\Gamma(\delta p)^{2}/2=q$, where $\Gamma$ is the slope of the latent liquidity. Once this is achieved, the price tends to revert back as $1/\sqrt{t-t_{i}}$, as predicted by the LLOB model, see [22, 7], but in disagreement with the predictions of the propagator model [16]. This discrepancy will be discussed further in [41].
If the above interpretation is correct, however, it should hold for arbitrary market orders. Since all orders are equivalent, they should lead to the same average impact on prices (as was indeed found in [31]). In other words, one should see a $\sqrt{q}$ impact for single market orders provided one waits long enough for the “hot potato” game to be completed. This is what we test in the next Section, which then opens up the question of reconstructing synthetic metaorders from a list of consecutive market orders that do not necessarily belong to the same ID. The anonymity of market orders suggests that, in certain conditions, the very same square-root impact law $\mathcal{I}$ given by Eq. (1) should also hold for synthetic metaorders. This indeed turns out to be the case, as we discuss now.
@@ 4.1 The impact of single public market orders
We want to understand the behavior of the price after a buy (sell) market order. Clearly, if the volume of the market order is less than the prevailing volume at the opposite best, the immediate impact is zero. However, as time goes by, one very quickly sees that impact grows and becomes approximately given by $\sqrt{q}$, whether or not immediate impact is zero (the full temporal aspects will be explored in more details in [41]).
@@ 4.2 Synthetic metaorders
Since arbitrary market orders seem to all behave similarly, independently of the metaorder they belong to, we made the following numerical experiment that confirms the non-linear propagator interpretation of Eq. (2) for metaorders. We construct a new dataset of synthetic metaorders, by randomly shuffling traders ID and distributing them to market orders while we keep the historical market order flow, i.e. by randomly reordering real traders’ IDs  [^3] The shuffling was based on the Fisher-Yates algorithm. In other words, we collect all the market orders within the same session for a specific stock and simply shuffle only trader IDs. Specifically, we apply the function numpy.random.shuffle in Python to the DataFrame column containing the trader IDs. This preserves the initial frequency distribution of traders: i.e. some of them appear many times whereas others are trading less frequently.
We then use the same method as in 2.1 to define metaorders as a consecutive sequence of trades of the same sign associated to the same new trader ID (that has been reshuffled). We obtain synthetic metaorders, that start and end at different times as the original (true) metaorders. Hence, any information associated with these metaorders is at least partially lost. Still, as shown in Fig. 6, we recover exactly the same square-root impact function as for the original metaorders! Note in passing that these graphs show once again that the square-root impact law only depends on the volume $Q$ of the metaorder (either real or synthetic) but not of the execution time $T$, see [9, 7, 25]. We have tested different constructions of synthetic metaorders on different stocks, from the Paris and London Stock Exchange, with similar results – a more detailed discussion will be presented in [41]. Note that the preservation of the impact law under reshuffling provides further evidence that short-term impact should be decoupled from alpha (i.e., predictive signals). This is consistent with the fact that the typical time horizon over which traders seek alpha is significantly longer than the execution timescale of metaorders. This is particularly true in this study, as metaorders are defined within trading periods. Similar observations have been made, for instance with CFM’s trades—where the square-root law remains valid—or in the ANcerno dataset, see [4, 6, 33]
Figure 6: Left: Impact $\mathcal{I}(Q)$ of real metaorders as a function of their rescaled size $Q/V_{D}$ for a typical Japanese Stock, using data from 2012 to 2018. The color of the dots correspond to different total execution time $T$, expressed in seconds. Right: Impact $\mathcal{I}(Q)$ of synthetic metaorders for the same stock, obtained via ID reshuffling. ID shuffling consists in a random permutation of historical trader IDs, preserving the frequency of apparition. Note that the vertical and horizontal scales are the same in the two plots: the square-root fit is exactly the same as for real metaorders. From the legend, one can also clearly see that $\mathcal{I}(Q)$ is independent of $T$ [9].
@@ 4.3 Discussion
The results of the previous two sub-sections strongly suggest a purely “mechanical” interpretation of the square-root impact law, based on a time decaying $\sqrt{q/t}$ impact of single market orders, independently of their association with a specific metaorder (since trader ID’s can be scrambled without affecting the results). These findings are difficult to reconcile with theories explaining the square-root law based on information, such as in [12], or on the detection by the market of the beginning of new metaorders, such as in [13, 21].
Those results are, on the other hand, perfectly in line with the fact that, due to anonymity, all market orders – even uninformed ones – should play an equivalent role and should on average impact prices similarly [20]. This was already noted in [31] by comparing the impact of CFM market orders with non-CFM market orders, and even more convincingly in an unpublished specifically designed 2010 experimental campaign with totally random market orders.
@@ 5.1 Refill sequences
Whereas the long-term correlation of market orders is well documented [16, 14, 17] and mainly attributed to the order splitting of large metaorders [14, 29], our dataset also allows us to study the splitting strategy of liquidity providers.
To do so, we simplify the problem by restricting to the set of filled limit orders, i.e. limit orders that have been placed in the order book and subsequently executed by another participant. Thus, as for liquidity takers, one can aggregate those filled limit orders into “refill sequences”, i.e. sequences of consecutive filled limit orders of same sign submitted by the same trader during a trading session. Given the splitting behavior of liquidity takers, market makers/liquidity providers face a sign-correlated succession of market orders. It is thus likely that the executed limit orders flow will be also be persistent, and lead to a power-law tail in the size distribution of the refill sequences, as we indeed confirm empirically, see Fig. 7. We show there a power-law fit of the distribution of the number $n$ of child orders associated with refill sequences, as $\psi(n)\propto n^{-\mu_{p}}$ with $\mu_{p}$ in the range [1.4,2.4][1.4,2.4][ 1.4 , 2.4 ] depending on the considered stock. This power-law decay echoes the Lillo-Mike-Farmer distribution of child market orders, although it is not expected to mirror it exactly since different traders provide liquidity to the same incoming metaorder – see [43] for a related discussion.
It should however be emphasized at this point that the boundary between liquidity provider and liquidity taker is somewhat blurred. Except for specific cases, the large majority of market participants use a mix of limit and market orders to acquire or sell shares. For example, it is quite common to see participants adding liquidity at the bid (ask) while sending market orders at the ask (bid). Large funds like AQR declare that most of their executed volume is though limit orders [36].
@@ 5.2 Strategic behaviour of liquidity providers
  • High $C_{\ell}\gtrsim 0.02$ traders are “wary” and place their next limit order quite a bit deeper in the book. The corresponding values of $\kappa_{\ell}$ cluster around 1/21/21 / 2.
@@
[…]
@@ 6 Conclusion
The JPX database provides a trove of interesting features, which have only started to be exploited by Sato and Kanazawa to understand the origin of the long memory of market order signs [29] and, more recently, to firmly establish the square-root law of market impact, Eq. (1) and rule out some of the proposed theories that predict a non-universal value for the concavity exponent $\delta$ [15].
@@
Our aim in this study was to leverage the fact that all metaorders can be identified to shed light on the microscopic origin of the square-root law. Our central result, which we did not expect when starting this project, is that such a law has in fact microscopic roots, and applies already at the level of single child orders, provided one waits long enough for the market to “digest” these orders. This is not consistent with the standard interpretation of the LLOB model [22], which assumed that the theory described the liquidity dynamics at a “mesoscopic” scale. The mesoscopic impact of metaorders rather arises from a “double” square-root effect at the level of child orders, see Eq. (2): square-root in volume of individual impact, followed by an inverse square root decay as a function of child order time, such that the cumulative impact of a metaorder yields back the partial (Eq. (4)) and total (Eq. (3)) square-root laws.
@@
This finding however immediately suggests that since market orders are anonymous, the double square-root law Eq. (2) should apply to any market orders and the impact of synthetic metaorders, reconstructed by scrambling the identity of the issuers, should also be described by the square-root impact law, Eq. (1). We have provided empirical evidence that this both statements are indeed valid. In particular, synthetic metaorders behave exactly as real metaorders, see Fig. 6. We conclude that there is nothing special about child orders belonging to a given metaorder, at odds with theories that emphasize the information content of such trades to explain the square-root impact law, but in agreement with previous conjectures about the purely mechanical aspect of price impact [20, 31, 11].
@@
Interestingly, our synthetic metaorder experiment suggests that it may be possible to reconstruct the impact of metaorders from the public tape only, without trader IDs. In a forthcoming publication [41], we show that this is indeed the case, provided market orders are properly aggregated into synthetic metaorders, opening the path to a new wave of empirical studies, in particular concerning cross-impact [44, 45].
@@
While our results show that the square-root impact law does not emerge at the meso-scale but is already present at the micro-scale, they trigger new unanswered questions. In particular, why is the impact of single market orders of volume $q$ also a square-root? We have argued that this is because the latent order book is locally linear, such that after a “hot potato” game between liquidity providers, the final counterparty of the initial market order is on average at a distance $\sqrt{q}$ from the initial mid-point. Although this scenario is intuitively plausible, we believe that a deeper dive into the JPX database (or a similar one) would allow one to (in-)validate such a picture. Another conundrum is the square-root time decay of individual market orders, which seems to violate the martingale constraint that relates the decay of the autocorrelation of the sign of market orders to impact decay [16, 40]. We hope to come back to these pressing issues in the near future.